### v2

# Turning Up the Heat: Min-p Sampling for Creative and Coherent LLM Outputs

Abstract
Large Language Models (LLMs) generate text by sampling the next token from a probability distribution over the vocabulary at each decoding step. However, popular sampling methods like top- (nucleus sampling) often struggle to balance quality and diversity, especially at higher temperatures, leading to incoherent or repetitive outputs. To address this challenge, we propose min- sampling, a dynamic truncation method that adjusts the sampling threshold based on the modelβs confidence by scaling according to the top tokenβs probability. We conduct extensive experiments on benchmarks including GPQA, GSM8K, and AlpacaEval Creative Writing, demonstrating that min- sampling improves both the quality and diversity of generated text, particularly at high temperatures. Moreover, human evaluations reveal a clear preference for min- sampling in terms of both text quality and diversity. Min- sampling has been adopted by multiple open-source LLM implementations, highlighting its practical utility and potential impact.
1 Introduction
Large Language Models (LLMs) have achieved remarkable success in generating coherent and creative text across diverse domains, from factual question answering to open-ended storytelling. A central challenge in these generative tasks is managing the trade-off between creativity and coherence, often influenced by the sampling strategy used during text generation. Popular methods like top- sampling (nucleus sampling) (Holtzman etΒ al., 2020) and temperature scaling (Ackley etΒ al., 1985) are widely adopted to address this challenge, but they often struggle, especially at higher temperatures. While increasing temperature can enhance diversity, it frequently reduces the coherence of generated text; conversely, more conservative sampling limits creativity and can lead to repetitive outputs.
This challenge becomes particularly critical when LLMs are used in tasks that require both imaginative and contextually grounded responses. In this paper, we address this fundamental issue by introducing a new sampling method called min- sampling, designed to dynamically balance creativity and coherence, even at high temperatures. Min- sampling establishes a minimum base probability threshold that scales according to the top tokenβs probability, allowing it to dynamically include diverse options when the model is uncertain while focusing on high-confidence tokens when the model is confident.
To demonstrate the effectiveness of min-, we conduct extensive experiments on various benchmark datasets, including GPQA (Rein etΒ al., 2023), GSM8K (Cobbe etΒ al., 2021), and AlpacaEval Creative Writing (Li etΒ al., 2023). Our results show that min- sampling outperforms top- and other popular decoding methods such as top- (Fan etΒ al., 2018) and -sampling (Hewitt etΒ al., 2022a), maintaining coherence while allowing for increased diversity, particularly with high-temperature scaling. We also conducted a comprehensive human evaluation to compare the quality and diversity of text generated by min- with that generated by traditional sampling methods. The results indicate a clear preference for min-, with participants rating min- outputs as superior in quality and diversity.
The contributions of this paper are as follows:
β’
We introduce min- sampling, a novel dynamic truncation method that effectively balances creativity and coherence in LLM-generated text, particularly at high temperatures.
β’
We present comprehensive experimental results on several benchmarks, demonstrating that min- consistently improves the quality and diversity of generated text compared to top- sampling and other existing methods.
β’
We validate the practical utility of min- through an extensive human evaluation, showing that human evaluators prefer min- outputs over those generated by other methods in terms of both quality and diversity.
β’
We provide practical empirical guidelines for using min- sampling, assisting practitioners in selecting appropriate hyperparameters and best practices for various applications.
β’
The rapid adoption of min- by the open-source LLM community further highlights its effectiveness and practical potential.
By introducing min- and offering empirical guidelines for its use, we aim to explore high-temperature settings for creative text generation without compromising coherence. Our results demonstrate that min- is a viable and superior alternative to existing sampling methods, both at standard and high-temperature settings, making it an important contribution to generative language modeling.
2 Related Work
Sampling methods are crucial in controlling the quality and diversity of text generated by LLMs. The choice of sampling strategy directly affects the balance between creativity and coherence, which is critical in many generative tasks. In this section, we review existing sampling methods and their limitations, establishing the motivation for our proposed min- sampling approach.
Greedy Decoding and Beam Search.
Greedy decoding and beam search are deterministic decoding strategies that select the token with the highest probability at each step (Freitag & Al-Onaizan, 2017). While these methods ensure high-probability token selection, they often lead to repetitive and generic text due to their lack of diversity. Beam search also incurs a significant runtime performance penalty.
Stochastic Sampling Methods.
Stochastic sampling methods aim to inject diversity into the generated text by introducing randomness in token selection. Temperature scaling adjusts the distributionβs sharpness, balancing diversity and coherence (Ackley etΒ al., 1985); however, higher temperatures often lead to incoherent and nonsensical results, limiting its applicability. Top- sampling selects from the top most probable tokens, ensuring that only high-probability tokens are considered (Fan etΒ al., 2018). While it offers a simple way to prevent unlikely tokens from being sampled, it does not adapt dynamically to varying confidence levels across different contexts.
Top- sampling, also known as nucleus sampling, restricts the token pool to those whose cumulative probability exceeds a predefined threshold (Holtzman etΒ al., 2020). This method effectively balances quality and diversity by focusing on the "nucleus" of high-probability tokens and dynamically adapts to different contexts. However, at higher temperatures, top- sampling can still allow low-probability tokens into the sampling pool, leading to incoherent outputs. This trade-off between creativity and coherence at high temperatures is a key limitation that we aim to address with min- sampling.
Entropy-Based Methods.
Recent work has introduced methods such as entropy-dependent truncation (-sampling) and mirostat sampling, which attempt to dynamically adjust the sampling pool based on the entropy of the token distribution (Hewitt etΒ al., 2022a; Basu etΒ al., 2021; ShwartzΒ Ziv & LeCun, 2024). While entropy/uncertainty-based approaches show promise in improving text quality, they often require complex parameter tuning and are computationally expensive, making them challenging to use in practical applications. We detail our experimental challenges running sampling in SectionΒ B.2.
3 Min- Sampling
The core idea of min- sampling is to dynamically adjust the sampling threshold based on the modelβs confidence at each decoding step. This dynamic mechanism allows the sampling process to be sensitive to the context and the certainty of the model, providing a better balance between creativity and coherence, especially at high temperatures.
3.1 Overview of Min- Sampling
In standard autoregressive generation, a language model predicts the probability distribution over the vocabulary for the next token, conditioned on the sequence generated so far. At each step, the model selects a token from this distribution either deterministically or stochastically. Min- sampling is a stochastic method that adapts its truncation threshold based on the modelβs confidence, allowing the sampling strategy to be context-sensitive.
Formally, at each time step , let denote the vocabulary, and represent the conditional probability distribution over the vocabulary for the next token . Min- sampling involves the following steps:
1.
Calculate the Maximum Probability: Identify the maximum probability token in the distribution, denoted as .
2.
Define the Truncation Threshold: Set a base probability threshold, , and scale it by to determine the actual truncation threshold:
(1) This threshold ensures that tokens with sufficiently high relative probabilities are considered while filtering out less probable tokens in a context-dependent manner.
3.
Define the Sampling Pool: Construct the sampling pool consisting of tokens whose probabilities are greater than or equal to :
(2) -
4.
Sample from the Pool: Sample the next token from the reduced set according to their normalized probabilities:
3.2 Intuition Behind Min- Sampling
The key intuition behind min- sampling is that token truncation thresholds are relative and depend on how certain the distribution is for that token, and not absolute thresholds. When the model is highly confident about the next token (i.e., is high), min- restricts the pool to high-probability candidates to maintain coherence. Conversely, when the model is less confident, relaxing the sampling pool allows for a more creative and diverse generation. Unlike top- sampling, which truncates the distribution based on a fixed cumulative probability, min- dynamically adjusts the threshold based on the modelβs confidence, leading to more context-sensitive generation.
Figure 1 illustrates the effects of different sampling methods, including min-, on token probability distributions. In subfigure (a), we show an initial probability distribution over tokens. Subfigures (b), (c), and (d) demonstrate how top-, top-, and min- sampling methods select tokens based on this distribution. Min- sampling dynamically adjusts its filtering threshold based on the modelβs confidence, focusing on high-probability tokens when confident and including diverse but plausible options when uncertain. This dynamic behavior helps min- balance coherence and diversity more effectively than top- and top- sampling.
3.3 Advantages Over Existing Methods
Min- sampling dynamically adjusts the sampling threshold based on the modelβs confidence, balancing creativity and coherence effectively. Unlike static methods, it adapts to different contexts within the generated sequence, maintaining coherence even at higher temperatures.
Balancing Creativity and Coherence.
Min- sampling effectively balances creativity and coherence by dynamically adjusting the sampling pool based on the modelβs confidence. In contrast, fixed thresholds used in methods like top- and top- sampling often lead to either overly diverse (and incoherent) or overly conservative (and repetitive) outputs. The dynamic nature of min- allows it to tailor its behavior to different contexts within the same generated sequence.
Robustness at High Temperatures.
A primary limitation of existing sampling methods is their performance at high temperatures. As the temperature increases, the token probabilities become more uniform, allowing unlikely tokens to be selected, which can result in incoherent text. Min- addresses this issue by scaling the truncation threshold proportionally to the modelβs confidence, ensuring that the output remains sensible even at higher temperatures. This capability is particularly valuable for tasks that benefit from high creativity, such as storytelling and dialogue generation.
Computational Efficiency.
Min- sampling retains computational simplicity, requiring only a few additional calculations over standard top- sampling. Unlike methods that involve auxiliary models or complex entropy-based adjustments, min- can be easily integrated into existing LLM inference pipelines without significant overhead. This makes it practical for both research and real-world applications, and offers a distinct advantage over other entropy-based methods such as and sampling (Hewitt etΒ al., 2022b), as we discuss in SectionΒ B.2.
3.4 Implementation Details
Implementing min- sampling requires minimal changes to standard language model decoding pipelines. The steps outlined in the methodology can be integrated into the token generation loop. Here are some practical considerations:
Integration into Decoding Pipelines.
Min- sampling can be implemented as a logits processor in frameworks like Hugging Face Transformers (Wolf etΒ al., 2020). After applying temperature scaling, the scaled threshold is computed, and tokens with probabilities below this threshold are filtered out before sampling. These operations are efficiently implemented using vectorized computations, adding negligible overhead to the decoding process.
Parameter Selection Guidelines.
β’
Choosing the Base Threshold (): Setting between 0.05 and 0.1 provides a good balance between creativity and coherence across various tasks and models. Higher values of (e.g., close to 1) can be used to maintain coherence at very high temperatures.
β’
Temperature Settings: Min- sampling works effectively across a wide range of temperatures. Practitioners can experiment with higher temperatures (e.g., or ) to enhance diversity without significant loss of coherence.
β’
Combining with Other Techniques: While min- sampling can be used in conjunction with other sampling methods or repetition penalties, it is recommended to use it as the primary truncation method to fully leverage its dynamic capabilities.
Ensuring Robustness.
To prevent the sampling pool from becoming empty, especially when is high and is low, it is advisable to enforce a minimum number of tokens to keep in .
3.5 Availability of Resources
To facilitate adoption, reference implementations of min- sampling are available:
β’
Hugging Face Transformers: An implementation is available as a custom logits processor that can be integrated into the generation pipeline.
- β’
β’
Project Repository: Code, usage examples, and integration guides are available at our project repository.111https://github.com/menhguin/minp_paper/
β’
Community Adoption: Min- sampling has been rapidly adopted by the open-source community, with over 54,000 GitHub repositories using it, amassing a cumulative 1.1 million stars across these projects.
This widespread community adoption highlights the practical utility and effectiveness of min- sampling in real-world applications. By following these guidelines and utilizing the available resources, developers can easily incorporate min- sampling into their language models to achieve an optimal balance between creativity and coherence with minimal effort.
4 Case Studies: Illustrative Examples
To provide qualitative insights into how min- sampling operates compared to existing methods, we present two case studies that highlight the differences in token selection, especially at higher temperatures. These examples illustrate the dynamic behavior of min- sampling in practice and set the stage for the comprehensive quantitative experiments that follow. This visualization was originally created by Maso (2024) and reproduced in this paper.
| (a) Case Study 1: Low-Certainty Next Token Prompt: "You will pay for what you have done," she hissed, her blade flashing in the moonlight. The battle that ensued ____ Token Top- Min- was 70.3 11.9 13.1 18.5 lasted 9.5 6.1 6.7 9.5 between 6.2 5.3 5.9 8.2 left 4.5 4.8 5.3 7.4 would 3.2 4.3 4.7 6.6 seemed 0.5 2.3 2.5 3.5 | (b) Case Study 2: High-Certainty Next Token Prompt: A rainbow is an optically brilliant meteorological event resulting from refraction, reflection, and dispersion of ____ Token Top- Min- light 98.3 34.4 38.2 80.9 sunlight 1.3 8.1 9.0 19.1 water 0.1 3.4 3.8 β sunshine 0.1 2.9 3.2 β a 0.05 2.7 3.0 β moisture 0.05 2.7 3.0 β |
Case Study 1: Low-Certainty Next Token
Prompt: "You will pay for what you have done," she hissed, her blade flashing in the moonlight. The battle that ensued _____
In this creative writing prompt, the model is expected to continue a story where multiple plausible continuations exist with multiple plausible continuations. The next token is uncertain, and the probability distribution is relatively flat at a high temperature.
Case Study 2: High-Certainty Next Token
Prompt: A rainbow is an optically brilliant meteorological event, resulting from the refraction, reflection, and dispersion of _____
In this factual prompt, "light" is the expected continuation, with the model highly confident in this token. We examine how various sampling methods manage this high-certainty context at .
Analysis and Insights
The case studies illustrate how min- sampling dynamically adjusts the sampling threshold based on the modelβs confidence, effectively balancing creativity and coherence. In low-certainty scenarios (Case Study 1), min- behaves similarly to top- sampling, allowing a range of plausible continuations and promoting diversity without sacrificing narrative coherence. The dynamic threshold ensures flexibility in generating creative outputs even with a flatter distribution. Conversely, in high-confidence scenarios (Case Study 2), min- prioritizes the most relevant tokens, effectively filtering out less pertinent options and maintaining factual accuracy and coherence even at high temperatures. This adaptability demonstrates min-βs ability to handle uncertain and confident contexts, ensuring robust performance by filtering out low-probability, potentially incoherent tokens.
5 Experiments
We comprehensively evaluated min- sampling compared to existing methods across multiple benchmarks and model sizes. Our experiments aimed to demonstrate that min- sampling effectively balances creativity and coherence, particularly at higher temperatures.
5.1 Experimental Setup
Models
Experiments were conducted using the Mistral 7B language model (Jiang etΒ al., 2023), selected for its strong performance across various tasks. To evaluate whether the benefits of min- sampling scale to larger models, we also performed tests on Mistral Large with 123B parameters.
Benchmarks
We evaluate min- sampling on three diverse benchmarks:
Sampling Methods and Hyperparameters
We compared min- sampling against baseline methods, including top- sampling (Holtzman etΒ al., 2020), temperature sampling, sampling (Hewitt etΒ al., 2022b), sampling (Hewitt etΒ al., 2022b) and mirostat sampling (Basu etΒ al., 2021). We present results between temperatures 0.7 and 3.0, with further tests between 0 and 5 linked in our project repository 222https://github.com/menhguin/minp_paper/.
For min-, base probability thresholds of and were used, while top- sampling employed . These hyperparameter settings were chosen based on empirical guidelines and prior research to provide a fair comparison (See SectionΒ B.1 for extensive discussion).
Evaluation Metrics
Evaluation metrics were tailored to each benchmark. For GPQA and GSM8K, we measured accuracy. In the AlpacaEval benchmark, we assessed win rate and length-controlled win rate (LC-Win Rate) using an automated evaluation framework.
5.2 Results
5.2.1 Graduate-Level Reasoning (GPQA Main)
Setup
The GPQA Main Benchmark consists of challenging, graduate-level multiple-choice questions in biology, physics, and chemistry. We used a 5-shot prompting strategy to provide context and improve performance, following standard practices (Rein etΒ al., 2023).
Results
TableΒ 2 presents the accuracy results on GPQA Main for different sampling methods and temperature settings using Mistral 7B. Min- sampling consistently achieves higher accuracy than others across all temperature settings. The performance gap widens at higher temperatures, demonstrating min-βs robustness in maintaining correctness even when increasing diversity.
Large Model Evaluation
We also evaluated min- sampling on the Mistral Large model with 123B parameters. The results, shown in TableΒ 3, indicate that the advantages of min- sampling persist with larger models, suggesting that its benefits scale with model size.
5.2.2 Grade School Math (GSM8K CoT)
Setup
The GSM8K CoT dataset comprises 8,500 grade school math word problems (Cobbe etΒ al., 2021). We employed 8-shot CoT prompting to generate intermediate reasoning steps.
Results
As shown in TableΒ 2, min- consistently outperforms the other methods in almost all temperature settings. The performance advantage becomes more pronounced at higher temperatures, indicating that min- sampling preserves the modelβs problem-solving abilities even when generating more diverse outputs. We also observed significant differences in test-time computing across sampling methods, as detailed in SectionΒ B.2, where and sampling exhibited exponential runtime increases with temperature compared to min-, and failed to load on altogether.
| Method | GPQA Main (5-shot) | GSM8K CoT (8-shot) | ||||||||
| Tempβ Only | 27.23 | 22.77 | 25.22 | 5.80 | 0.89 | 29.56 | 17.51 | 0.00 | 0.00 | 0.00 |
| Top- | 26.34 | 23.66 | 22.77 | 16.52 | 5.88 | 30.63 | 17.59 | 0.00 | 0.00 | 0.00 |
| Sampling | 28.13 | 25.45 | 24.55 | β | β | 32.63 | 26.99 | 0.49 | β | β |
| Sampling | 27.90 | 25.45 | 24.11 | β | β | 31.69 | 26.84 | 0.56 | β | β |
| Top- | 29.02 | 25.00 | 24.78 | 6.47 | 0.46 | 36.09 | 27.67 | 0.68 | 0.00 | 0.00 |
| Min- | 29.18 | 25.89 | 28.13 | 26.34 | 24.55 | 35.18 | 30.86 | 18.42 | 6.21 | 0.00 |
Accuracy vs. Diversity Trade-off
To further understand the accuracy-diversity tradeoff, we evaluate both metrics on the GSM8K dataset using chain-of-thought reasoning with using self-consistency decoding (Wang etΒ al., 2022). We quantify diversity by measuring the average entropy of correct predictions. Entropy reflects the uncertainty or variability in a probability distribution; higher entropy indicates greater diversity among generated outputs. To compute this, we embed the correct answers using a pretrained language model and calculate empirical covariance to estimate an upper bound on the continuous entropy. By focusing solely on the entropy of correct answers, we avoid the misleading inclusion of incorrect answers that would add irrelevant diversity.
The results shown in FigureΒ 2 illustrate that min- sampling achieves a better trade-off between accuracy and creativity compared to top- sampling. Min- sampling consistently lies closer to the Pareto frontier, indicating more efficient performance. The greater spread of min- configurations shows its sensitivity to hyperparameter settings, allowing fine-grained control over the diversity and coherence of outputs, whereas top- configurations cluster strongly, showing that top- sampling is less sensitive to hyperparameter values. We further discuss this nuance in SectionΒ B.3.
5.2.3 Creative Writing
Setup
We used the AlpacaEval Creative Writing benchmark to assess the modelβs ability to generate creative and engaging text (Li etΒ al., 2023). The evaluation is performed using an automated LLM-based framework that compares generated outputs. Similarly to Gusev (2023), we report both the win rate and the length-controlled win rate (LC-Win Rate), which controls for differences in output length.
Results
TableΒ 3 shows that min- sampling outperforms both top- sampling, sampling and Mirostat. Min- achieves a significantly higher win rate, indicating its effectiveness in producing high-quality creative writing without sacrificing coherence.
5.3 Ablation Study
We conducted an ablation study on the AlpacaEval Creative Writing benchmark to evaluate the impact of different parameter settings on the performance of the min- sampling method. Specifically, we compared min- and top- sampling across various temperatures and configurations. We used two key metrics: Winrate and Winrate (LC).
The results in Table 6 in the Appendix show that min- sampling generally outperforms top- sampling across different temperatures and parameter settings. The highest winrate is achieved with at temperature , demonstrating that min- sampling is effective in producing high-quality outputs even under conditions that promote creativity (higher temperatures). Moreover, the Winrate (LC) results are consistent with the Winrate, confirming that the benefits of min- sampling are robust to biases due to differences in output length.
6 Human Evaluation
To complement our quantitative benchmarks and explore the qualitative benefits of min- sampling, we conducted a comprehensive human evaluation focusing on creative writing. This evaluation aimed to assess the perceived quality and diversity of text generated using min- and top- sampling at various temperature settings.
Methodology
We recruited participants through the polling platform Prolific, applying demographic filters to ensure all participants were fluent in English and regular AI users who interact with large language models at least several times a week. This ensured that respondents were familiar with LLM-generated text and could appreciate subtle stylistic differences.
To ensure high-quality responses, we implemented several measures:
1.
Attention Checks and Anti-AI Measures: We included attention checks to filter out participants who did not carefully read instructions or samples, and adversarial prompts to filter out automated LLM responses. Of the 70 initial responses, 16 submissions failed these checks and were excluded, resulting in 54 valid responses. We provide an example of our attention check for reference:
2.
Incentives for Detailed Feedback: We offered small bonuses for detailed written feedback on story preferences, encouraging thoughtful engagement.
| (a) Accuracy (%) on GPQA Main benchmark (Mistral Large) Method Tempβ Only 37.72 31.25 29.02 20.09 2.90 0.89 Top- 39.51 33.26 29.24 18.75 2.01 0.67 Top- 40.18 34.38 29.69 21.21 2.01 0.89 Min- (ours) 38.17 34.60 31.03 27.46 22.77 13.84 | (b) Win rates (%) on AlpacaEval Creative Writing benchmark Method Temperature Only 49.97 53.18 Mirostat 16.69 14.23 Sampling 43.50 45.51 Top- 50.07 β Min- (ours) 52.01 56.54 |
Experimental Setup
For this creative writing task, we prompted a Llama 3 70B model with the prompt "Write me a creative story?" at various top-, min- and temperature settings. For each sampling method, outputs were generated at three temperature settings: , , and . For each temperature, we tested two diversity levels: low diversity with top- = 0.1 and min- = 0.2, and high diversity with top- = 0.9 and min- = 0.05. Participants were presented with three samples for each configuration to assess output diversity.
Evaluation Criteria
Participants rated each set of outputs on two criteria, using a scale from 1 (lowest) to 10 (highest):
1.
Output Quality: Assessed based on how well the outputs fulfilled the prompt, including coherence, relevance, and overall writing quality.
2.
Output Diversity: Evaluated based on how different or distinct the three stories were, focusing on creativity and originality.
Results
TableΒ 4 summarizes the average scores for quality and diversity across different temperature and diversity settings. Overall, min- sampling consistently scored higher than top- sampling across all settings. At higher temperatures, while top- samplingβs scores for quality and diversity decreased significantly, min- sampling maintained high scores. A paired t-test confirmed that the differences in scores between min- and top- sampling were statistically significant (p < 0.05).
Qualitative Feedback
Participants frequently noted that outputs generated with min- sampling were more coherent and creative, especially at higher temperatures. In contrast, top- sampling often produced incoherent, less diverse outputs in similar conditions, especially with low diversity settings.
| Temperature | Diversity Setting | Min- | Top- | ||
| Quality | Diversity | Quality | Diversity | ||
| 1.0 | Low | 7.06 1.48 | 5.83 2.03 | 5.96 2.24 | 2.40 2.01 |
| High | 8.02 1.35 | 7.74 1.63 | 7.67 1.38 | 7.04 1.88 | |
| 2.0 | Low | 7.62 1.53 | 6.91 1.94 | 5.43 2.24 | 1.83 1.61 |
| High | 7.98 1.42 | 7.96 1.54 | 7.75 1.37 | 7.66 1.50 | |
| 3.0 | Low | 7.74 1.76 | 7.60 1.86 | 5.75 2.33 | 2.25 2.44 |
| High | 7.57 1.68 | 7.66 1.45 | 7.11 2.09 | 7.49 1.74 | |
These results demonstrate that min- sampling is better in both output quality and diversity.
7 Conclusion
In this paper, we introduced min- sampling, a novel truncation sampling method for large language models that dynamically adjusts the sampling threshold based on the modelβs confidence at each decoding step. Our approach effectively balances creativity and coherence, particularly at higher temperatures where traditional methods like top- sampling often struggle.
Through comprehensive experiments across diverse benchmarks, we demonstrated that min- sampling consistently outperforms existing methods in both quality and diversity of outputs. Extensive human evaluations further confirmed a strong preference for min- sampling over top-, highlighting its practical advantages in real-world applications.
The key strengths of min- sampling are its simplicity, computational efficiency, and ease of integration into existing pipelines. By enabling models to generate text that is both creative and coherent, min- sampling addresses the longstanding trade-off between diversity and quality in text generation.
Min- sampling represents a significant advancement in generative language modeling, potentially enhancing a wide range of applications requiring high-quality and diverse text generation.
Reproducibility Statement
We have made significant efforts to ensure the reproducibility of our results. The implementation of the proposed min- sampling method is provided in AppendixΒ B and is also available anonymously at our project repository.333hhttps://github.com/menhguin/minp_paper/ Detailed descriptions of experimental setups, including model configurations, hyperparameter settings, and evaluation protocols, are outlined in SectionΒ 5 and SectionΒ B.1. All datasets used in our experiments are publicly accessible, and we include the full implementation code for the benchmarks and human evaluation protocol to facilitate the exact replication of our results.
Ethics Statement
Min- sampling aims to improve the diversity and coherence of text generated by large language models. We acknowledge the following ethical considerations:
β’
Potential misuse: Min- could potentially enhance the fluency of misleading or harmful content. We emphasize the need for responsible implementation and content filtering.
β’
Safety risks: It is possible that high-temperature text generation increases risks of circumventing safety finetuning, although, in practice, we are not aware of such instances.
β’
Transparency: To ensure reproducibility and further research, we have open-sourced our implementation and provided extensive details on the experimental setup and results. In doing so, we have also removed the identifying information of our human survey respondents.
We believe the benefits of entropy and uncertainty-based methods outweigh these risks. We strongly encourage safety and alignment research leveraging uncertainty and entropy, as this can significantly benefit robustness, truthfulness and reduced hallucinations (Stolfo etΒ al., 2024; Wang & Zhou, 2024).
Acknowledgements
Min- began as a passion project to see if small, open-source language models could write diverse and interesting stories. As such, its creation is a winding story with a diverse set of characters.
We thank Apart Research, particularly Esben Kran and Jason Schreiber, for providing extensive guidance throughout our Apart Lab Fellowship, during which min- evolved from an interesting idea to a full, empirically validated research paper. We also thank Wand AI for their generous sponsorship and assistance with evaluations on Mistral Large and Llama3 70B.
Our gratitude extends to the numerous members of the open-source LLM community who supported the research and adoption of min-. We are particularly indebted to Ilya Gusev for his independent evaluations, Romain Dal Maso for his intuitive visualizations, Hailey Schoelkopf for her extensive technical support, and JoΓ£o Gante for his help with the Hugging Face Transformers library.
While our coauthors are employed in various research roles, min- does not belong to any institution. Our technical contributions and research findings are free for use by the open-source community.
Contribution Statement
AB conceived the original idea of min- sampling, tested and shared it with hobbyists within the open-source LLM community, leading to its rapid and widespread external adoption.
MN initiated, planned and wrote most of the draft and final versions, conducted extensive GPQA and GSM8K evaluations on Mistral 7B, and coordinated with coauthors, collaborators and inference providers throughout the research process.
CN acted as research supervisor under the Apart Research Fellowship, oversaw initial drafts, handled paper formatting, early revisions and data visualization.
AK conducted GSM8K Self-Consistency evaluations to assess Diversity-Accuracy tradeoffs and helped flesh out theoretical explanations of min-.
AR conducted evaluations on Mistral Large and Llama 3 70B, coordinated research collaboration with Wand AI, generated outputs for human evaluation, and provided insights on optimal settings.
RSZ rewrote a substantial portion of the final draft, proposed the inclusion of a human evaluation, and provided guidance on evaluations, data presentation and academic publishing standards.
Appendix A Limitations and Future Work
While min- sampling shows significant promise, there are limitations and opportunities for future research:
Generalization to Other Models:
Our experiments focused on the Mistral 7B and Mistral Large language models. Future work should explore the effectiveness of min- sampling with larger models and different architectures to assess its generalizability.
Hyperparameter Sensitivity:
The base probability threshold is a critical hyperparameter. Investigating methods for dynamically adjusting based on context or developing guidelines for optimal settings across various tasks could enhance performance. This was particularly challenging, as MAUVE hyperparameter sweeps are measured without temperature scaling on GPT2-XL (Pillutla etΒ al., 2023), and require pre-selection on exact hyperparameters. min-, however, is a novel method often used in conjunction with temperature scaling without extensive literature/experimental data. We discuss this further in SectionΒ B.1.
Theoretical Analysis:
A deeper theoretical understanding of why min- sampling performs better, particularly at high temperatures, could provide insights into the behavior of language models and guide the development of even more effective sampling strategies. A possible direction can be using information theory (Shwartz-Ziv etΒ al., 2023)
Applicability to Other Domains:
Extending min- to other generative tasks, such as code generation or multimodal models, could reveal broader applicability and benefits across different domains.
Research into high-temperature regimes:
High-temperature regimes have been underexplored relative to low-temperature regimes. Min- sampling hopes to unlock exploration, experimentation, and applications in such areas.
Human Evaluation Scope:
Our human evaluation involved participants selecting pre-generated outputs. We note that min-βs popularity within the open source community for creative writing is interactive in nature; hence, we hope for adoption on interactive human evaluation platforms such as the Chatbot Arena (Chiang etΒ al., 2024).
Combining uncertainty and CoT decoding methods:
Wang & Zhou (2024) found a significant correlation between the confidence/certainly level of the final answer token choice and correct scores on GSM8K CoT, and that promoting diverse lower-probability token choices encouraged generating chains of thought that were beneficial for reasoning, resulting in higher scores overall.
This mirrors our hypothesis that choosing high-certainty tokens enables more accurate final answers, while diverse token choices benefit intermediate reasoning steps. For example, we note that GPQA scores on Mistral 7B increased from to , but only with Temperature Only and min-. With the recent release of OpenAIβs O1 models, which leverage CoT methods at inference for advanced reasoning capabilities, we note several novel decoding methods that combine uncertainty and CoT sampling approaches to improve model reasoning in a simple manner, with minimal added overhead and architectural changes (Wang & Zhou, 2024; Shrek & Frog, (2024) . We aim to actively explore such methods in future work.
Appendix B Min- Implementation and Documentation
Below is the implementation code for min- truncation sampling as detailed in the Hugging Face Transformers library, with range exception handling and keeping minimum tokens to prevent errors.
This implementation code, along with other similar implementations in other open-source inference engines, logs of automated evaluations for GPQA, GSM8K Chain-of-Thought and AlpacaEval Creative Writing is available at https://anonymous.4open.science/r/minp_paper-767F/.
B.1 Hyperparameters Settings
To choose fair and optimal hyperparameter settings, we mainly cross-referenced publicly-reported scores on MAUVE (Pillutla etΒ al., 2023), common recommendations from leading model providers, and any recommendations from the original authors. We also found that Risk Levels (Zhou etΒ al., 2024) correlate strongly with optimal results across temperature ranges. Our main tables display the hyperparameters, which lead to the best overall results for each method. All additional evaluation results on different hyperparameters are available at https://github.com/menhguin/minp_paper/
For min-, base probability thresholds of and were used, while top- sampling employed .
For min- = 0.05 and 0.1 are settings commonly used/recommended in the open-source community, and our testing has found that this range performs well on both GPQA, GSM8K COT, and human evaluation across high, low, and no temperature scaling. We also tested min- = 0.2 and min- = 0.3, but these are not commonly used.
For top-, top- = 0.9 and top- =0.95 are settings commonly used/recommended in the open-source community, and found in several independent MAUVE assessments to be optimal (Hewitt etΒ al., 2022a; Zhu etΒ al., 2024). We mainly reference the Risk Levels framework from Zhou etΒ al. (2024), which measures tradeoffs between diversity and risk/precision in text generation, specifically Risk Level 15 for Mixtral 7B, which we used as a reference point for the top-, and sampling settings.
| Model | Method | Parameter | Risk Std Error β | Recall β |
| Mixtral-7b | Top-k | 181 | 1.759 | 0.364 |
| Top-p | 0.9315 | 6.315 | 0.447 | |
| Adaptive | 2.2e-5 | 2.757 | 0.466 | |
| Eta | 1.96e-4 | 4.712 | 0.505 | |
| Mirostat | 6.71 | 2.213 | 0.468 |
For top-, we could not find clear recommendations on the optimal hyperparameters. We conducted tests on k = 10, 15, 20, 40, 50 and 180. Due to the nature of top-, we noted that best scores and settings varied significantly by temperature, making a fair comparison difficult as, in practice, top- is meant to be a static threshold and not dynamically adjusted at inference. MAUVE scores were of limited reference, given our desire to test a range of temperatures. Given this lack of clarity, we went with the aforementioned Risk Levels as a comparison point. (Zhou etΒ al., 2024)
For and , we found inter-agreement between the authorβs original recommendation, independent MAUVE assessments (Zhu etΒ al., 2024), and Risk Levels. We tested and values 0.0002 and 0.0009, found 0.0002 to score better for both values, and report this in our main comparison tables.
B.2 Test Time Compute challenges
While running GPQA and GSM8K CoT for and sampling via Hugging Face and the EleutherAI Evaluation Harness (Gao etΒ al., 2023), we noted that test-time compute increased exponentially with temperature. Throughout our experiments, min-, top-, and top- generally took 2-5 minutes to evaluate on Mistral 7B at every temperature for both GPQA and GSM8K. Runtime on an A100 Colab increased from Β 5 minutes at to 8 minutes at and 30 minutes at . Neither nor seemed to function at . On GSM8K, and each took 2 hours to evaluate, which is equal to the average for our Llama 70B/Mistral Large run. This time also appeared to scale exponentially with increased temperature.
Our experience suggests min- is a practical alternative to and samplingβs entropy-based heuristics both on quantitative benchmarks and compute efficiency.
B.3 How percentages thresholds differ for min- and top-
In choosing hyperparameter values, min- and top-βs percentage thresholds differ in subtle but meaningful ways. Strictly speaking, min-βs threshold is not the same as an equivalent "top-- 1" threshold. For example, when top- = 0.9, the last <10% of the total distribution is truncated. However, itβs possible for min- = 0.1 to truncate more than 10% of a distribution.
Consider the following top 5 tokens probabilities: 80%, 7%, 3%, 2%, 1%. With top- set to 0.9, the top 3 tokens comprising 90% of the distribution is preserved. With min- = 0.1, and the resulting truncation threshold at 8%, only the top token is preserved, and 20% of the original distribution is truncated.
In practice, this means that in high-certainty token distributions, min- truncates a larger percentage of that probability distribution than its value. This contributes to min-βs ability to consistently choose high-certainty tokens despite high temperature scaling.
Hence, low values (such as from 0.01 to 0.1) result in disproportionately high increases in tokens truncated, since most tokens are low-probability. This results in Figure Β 2βs observation that min-βs is more sensitive than top-βs when adjusted by the same percentage values/basis points.
Appendix C Additional Results
C.1 Ablation Study
The results in Table 6 show that min- sampling generally outperforms top- sampling across different temperatures and parameter settings, particularly in terms of the Winrate metric. The highest winrate is achieved with at temperature , demonstrating that min- sampling is effective in producing high-quality outputs even under conditions that promote creativity (higher temperatures). Moreover, the Winrate (LC) results are consistent with the Winrate results, confirming that the benefits of min- sampling are robust to biases due to differences in output length.
| Method | Temperature | Configuration | Winrate (%) | Winrate (LC) (%) |
| Top- Sampling Configurations | ||||
| Top- | 0.8 | 54.65 | 51.29 | |
| Top- | 1.0 | 53.00 | 50.43 | |
| Top- | 1.0 | 52.07 | 50.07 | |
| Top- | 0.8 | 51.80 | 50.22 | |
| Top- | 0.8 | 50.76 | 48.78 | |
| Min- Sampling Configurations | ||||
| Min- | 1.5 | 58.12 | 56.54 | |
| Min- | 1.0 | 55.07 | 52.01 | |
| Min- | 1.0 | 53.24 | 50.14 | |
| Min- | 1.0 | 51.62 | 50.43 | |
| Min- | 1.0 | 51.46 | 48.85 | |
| Min- | 0.8 | 50.99 | 47.84 | |
C.2 GPQA
These results demonstrate min-βs ability to maintain higher levels of coherence and accuracy in multi-step reasoning tasks, even when the diversity of token selection is increased through higher temperature settings. This finding aligns with our hypothesis that min- sampling can better navigate the creativity-coherence tradeoff in complex reasoning scenarios.
C.3 Full Plots for GSM8K/GPQA results